OpenStudy (anonymous):

The area of circle with radius 'r' is the same as the area of a square of side 'a'. Ratio of 'r' to 'a' is:

OpenStudy (raden):

area of a circle = area of a square pi r^2 = a^2 r^2/a^2 = 1/pi (r/a)^2 = 1/pi r/a = sqrt(1/pi) = sqrt(1)/sqrt(pi) = 1/sqrt(pi)

OpenStudy (raden):

if you want rationalize it, times by sqrt(pi)/sqrt(pi). get r/a = 1/sqrt(pi) * sqrt(pi)/sqrt(pi) r/a = sqrt(pi)/(pi)